Ask your own question, for FREE!
Mathematics 8 Online
OpenStudy (anonymous):

An experiment consists of rolling a fair die and flipping a fair coin. What is the probability of getting an even number and a tail? 1/2 correct?

OpenStudy (shubhamsrg):

3/6 * 1/2

OpenStudy (shubhamsrg):

=1/4

OpenStudy (chaise):

3/6 * 1/2 = 1/6 There is a three in six chance of getting a even number on a dice with numbers 1-6, and a 1 in 2 chance of getting a tail on a coin. Multiply the two chances together.

OpenStudy (anonymous):

3/6 from where??

OpenStudy (anonymous):

nvm thx

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
1 day ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!